OpenStudy (anonymous):

Travis is buying bunny ears which cost $4.00 each. He also has a $2.00 off coupon. The equation y = 4x - 2 gives his final cost, y, to purchase x pairs of bunny ears. If x = 6, what is the ordered pair (x, y)?

OpenStudy (kirbykirby):

y = 4x - 2... now if x=6, plug it in: y=4(6) - 2 y=24-2 y=22 The ordered pair is (6, 22)